Answer the following problems.


2.) Determine the sum pf the real zeros of P(x) = 3x4 + 14x3 + 10x2 - 14x + 3.


3.) Find μ so that x - 3 is a factor of μx3 - 13x2 - μx - 3.



1
Expert's answer
2021-01-29T16:56:18-0500

2) The sum of the real zeros of P(x) is


"-\frac{b}{a}=-\frac{14}{3}"

3) By the Factor Theorem, x-3 is a factor of 

"f(x)=μx^3 - 13x^2 - μx - 3\\f(3)=3^3μ - 13(3)^2 -3 μ - 3=0\\\mu=5"